Expand the triple product for "\hvec{a}=\omega\times(\omega\timesr)." . if r is perpendicular to ω , show that "\hvec{a}=-\omega^2\hvec{r}"  and so find the elementary result that the acceleration is toward the center of the circle of magnitude "\frac{\upsilon^2}{r}"


1
Expert's answer
2021-03-01T11:50:28-0500
"\bold{\omega \times(\omega \times r)=(\omega\cdot r)\omega-(\omega\cdot\omega)r}"

"(\omega\cdot r)=0"

"\bold{\omega \times(\omega \times r)=-\omega^2r}"

"\bold{a=-\omega^2r=-e_r}\frac{v^2}{r}"